Evaluate the limits

"Lim"

Xto + sum to infinity


5xcube-2xsquared +1 divide by 1-3x


1
Expert's answer
2021-05-07T10:23:29-0400

We need to find,

"\lim_{x \to \infty} \frac{5x^3-2x^2}{1-3x}"


The highest power of the x is 3.

When we find "\lim_{x \to \infty}" we need to find the ratio of the coefficients of the highest power in numerator and denominator.

Here, "\lim_{x \to \infty} \frac{5x^3-2x^2}{1-3x} = \frac{5}{0 } = \infty"